The JED Editor 1

Contents

1 Introduction 2 2 Installing JED 2 3 Startup Files 3 4 Starting JED 3 5 Emulating Other Editors 4

5.1 Emacs Emulation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 5.2 EDT Emulation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 5.3 Wordstar Emulation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5

6 File Types and Sizes 6 7 Backup and Auto-save Files 6 8 Status line and Windows 6 9 Mini-Buffer 7

9.1 Command Line Completion . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 9.2 File Names . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 9.3 Buffer Name and File Name Completion . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9

10 Basic Editing 9 10.1 Undo . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 10.2 Marking Text (Point and Mark) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 The JED Editor 2

10.3 Tab Issues. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 10.4 Searching . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 10.5 Rectangles . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 10.6 Sorting . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14

11 Modes 15

11.1 Wrap Mode . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16

11.1.1 Formatting paragraphs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 11.2 Smart Quotes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 11.3 C Mode . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 11.4 Fortran Mode . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17

12 Keyboard Macros 18

13 Shells and Shell Commands 19

14 Getting Help 19

15 Editing Binary Files 19

16 Dired— the Directory editor 20

17 Mail 20

18 Customization 21 18.1 Setting Keys . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 18.2 Predefined Variables . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 18.3 Hooks . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 The JED Editor 3

18.4 S-Lang Programming Hints (Debugging) . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25

19 Eight Bit Clean Issues 27 19.1 Displaying Characters with the High Bit Set . . . . . . . . . . . . . . . . . . . . . . . . . . 27 19.2 Inputting Characters with the hight bit Set . . . . . . . . . . . . . . . . . . . . . . . . . . 27 19.3 Upper Case - Lower Case Conversions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28

20 Miscellaneous 29 20.1 Abort Character . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 20.2 Input Translation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 20.3 Display Sizes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30

1 Introduction

This document presents some basic information that users should know in order to use jed effectively. Any questions, comments, or bug reports, should be email-ed to the author. Please be sure to include the version number. To be notified of future releases of jed, email to the address below and your email address will be placed on the announcement list. The email address is:

      davis@space.mit.edu  (Internet)
      davis@ohstpy (BITNET)

2 Installing JED

Building jed from its sources requires the use of a C compiler that understands function prototypes. jed has been successfully built with cc on the ULTRIX, VMS, and IRIX operating systems. In addition, it has been created using gcc under SunOS and Borland’s BCC 3.0 for the PC.

Detailed installation instructions are in separate, operating system dependent files. They are:

       UNIX:   install.unx
        VMS:   install.vms
      IBM-PC:   install.pc

When jed starts up, it will first try to load a site initialization file called site.sl. Site specific commands are placed here. Most likely, site.sl will define some functions, default hooks, etc… What goes in it is left to the discretion of the user or system manager. See the file site.sl for examples.

When loading site.sl as well as other S-Lang files (the user’s personal initialization file, .jedrc or jed.rc, is a special case, see below), jed searches all directories specified by the environment variable JED_LIBRARY, and if the file is not found, jed will look for it in the default directory. The environment variable JED_LIBRARY is a comma separated list of directories. Here are some examples of setting this variable for different systems:

      VMS:    define/job JED_LIBRARY  dev$lib:[jedfiles]
      UNIX:   setenv JED_LIBRARY '/usr/local/lib/jed,~/jed'
      IBMPC:  set JED_LIBRARY = c:\editors\jed\lib

You will probably want to put define JED_LIBRARY in your login startup file, e.g., autoexec.bat,

login.com, or .cshrc.

jed versions 0.92 and later allow the value of JED_LIBRARY to be specified at compile time an it may only be necessary to define JED_LIBRARY as an environment variable to override its pre–compiled value.

3 Startup Files

Once jed has loaded the startup file site.sl, it will try to load the user’s personal initialization file. It first looks in the directory pointed to by the environment variable JED_HOME. If that fails, it then searches the HOME directory and upon failure simply loads the one supplied in JED_LIBRARY.

The name of the user initialization file varies according to the operating system. On Unix systems this file must be called .jedrc while on VMS and MSDOS, it goes by the name jed.rc. For VMS systems, the HOME directory corresponds to the SYS$LOGIN logical name while for the other two systems, it corresponds to the HOME environment variable.

The purpose of this file is to allow an individual user to tailor jed to his or her personal taste. Most likely, this will involve choosing an initial set of key-bindings, setting some variables, and so on.

4 Starting JED

Normally, jed is started as

     jed <file name>

or

     jed <file list>

However, jed also takes the switches defined in the following table:

     Switch          Meaning
     ------          -------
      -batch         run JED in batch mode.  This is a non-interactive mode.
      -n             do not load jed.rc (.jedrc) file.
      -g <n>         goto line `n' in buffer
      -l <file>      load `file' as S-Lang code.
      -f <function>  execute S-Lang function named `function'
      -s <string>    search forward for `string'
      -2             split window
      -i <file>      insert <file> into current buffer.

For example, the command line:

                jed slang.c -g 1012 -s error -2 file.c -f eob

will start up jed, read in the file slang.c, goto line 1012 of slang.c and start searching for the string error, split the window, read in file.c and goto the end of the file.

If the -batch parameter is used, it must be the first parameter. Similarly, if -n is used, it must also be the first parameter unless used with the -batch parameter in which case it must the second. jed should only be run in batch mode when non–interactive operation is desired. For example, jed is distributed with a file, mkdoc.sl, that contains S–Lang code to produce a help file for functions and variables. In fact, the help file jed_funs.hlp was created by entering

                jed -batch -n -l mkdoc.sl

at the command line.

Now suppose that you want to read in a file with the name of one of the switches, say -2. How can this be done? The answer depends upon the operating system. For Unix, instead of jed -2, use jed ./-2; for VMS, use jed []-2. The case for MS-DOS is similar to Unix except that one must use the backslash.

It is possible to completely change jed’s command line syntax through the use of the user defined function command_line_hook. In fact, the behavior described above is dictated by the value of command_line_hook as distributed in site.sl. See the section on hooks for details.

5 Emulating Other Editors

jed’s ability to create new functions using the S–Lang programming language as well as allowing the user to choose key bindings, makes the emulation of other editors possible. Currently, jed provides reasonable emulation of the Emacs, EDT, and Wordstar editors.

5.1 Emacs Emulation

Emacs Emulation is provided by the S-Lang code in emacs.sl. The basic functionality of Emacs is emulated; most Emacs users should have no problem with jed. To enable Emacs emulation in jed, make sure that the line

      evalfile ("emacs"); pop ():

is in your jed.rc (.jedrc) startup file. jed is distributed with this line already present in the default

jed.rc file.

5.2 EDT Emulation

For EDT emulation, edt.sl must be loaded. This is accomplished by ensuring that the line

      evalfile ("edt"); pop ();

is in present in the jed.rc (.jedrc) Startup File. jed is distributed with EDT emulation enabled on VMS and Unix systems but the above line is commented out in the jed.rc file on MS-DOS systems.

This emulation provides a near identical emulation of the EDT keypad key commands. In addition, the smaller keypad on the newer DEC terminals is also setup. It is possible to have both EDT and Emacs emulation at the same time. The only restriction is that emacs.sl must be loaded before edt.sl is loaded.

One minor difference between jed’s EDT emulation and the real EDT concerns the Ctrl-H key. EDT normally binds this to move the cursor to the beginning of the line. However, jed uses it as a help key. Nevertheless, it is possible to re-bind it. See the section on re-binding keys as well as the file edt.sl for hints. Alternatively, simply put

      unsetkey ("^H"); setkey ("bol", "^H");

in the jed.rc startup file after edt.sl is loaded. Keep in mind that the Ctrl-H key will no longer function as a help key if this is done.

EDT emulation for PCs only work with the enhanced keyboard. When edt.sl is loaded, a variable NUMLOCK_IS_GOLD is set which instructs jed to interpret the Num-Lock key on the square numeric keypad to function as the EDT GOLD key. In fact, this keypad should behave exactly like the keypad on VTxxx terminals. The only other problem that remains concerns the + key on the PC keypad. This key occupies two VTxxx key positions, the minus and the comma (delete word and character) keys. Thus a decision had to be made about which key to emulate. I chose the + key to return the characters Esc O l which jed maps to the delete character function. This may be changed to the delete word function if you prefer. See the file edt.sl for details.

The GOLD–GOLD key combination toggles the keypad between application and numeric states. On the PC, this is not possible. Instead, the PC F1 key has been instructed to perform this task.

5.3 Wordstar Emulation

wordstar.sl contains the S-Lang code for jed’s Wordstar emulation. Adding the line

      evalfile ("wordstar"); pop ();

to your jed.rc (.jedrc) startup file will enable jed’s Wordstar emulation.

6 File Types and Sizes

jed is primarily a text editor; however, it can also edit binary files (see the section on editing binary files). As a result, jed may edit lines of arbitrary length (actually this depends upon the size of an integer). It is capable of editing arbitrarily large buffers as long as there is enough memory for the buffer as well as the overhead involved. This editor employs a linked list representation; hence, the overhead can be quite high.

7 Backup and Auto-save Files

On UNIX and MS-DOS systems, jed creates backup files by appending a character to the filename. The VMS operating system handles backup files itself. jed periodically auto-saves its buffers. On UNIX and MS-DOS, auto-save files are prefixed with the pound sign #. On VMS, they are prefixed with \_$. The auto-save interval may be changed by setting the variable MAX_HITS to the desired value. The default is 300 “hits” on the buffer. A “hit” is defined as a key which MAY change the state of the buffer. Cursor movement keys do not cause hits on the buffer.

Like many of jed’s features, the names of auto-save and backup files can be controlled by the user. The

file site.sl defines two functions, make_backup_filename, and make_autosave_filename that generate

the file names described in the previous paragraph. Like all S–Lang functions, these functions may be overloaded and replaced with different ones. See also information about find_file_hook in the section on hooks.

On UNIX systems, jed catches most signals and tries to auto-save its buffers in the event of a crash or if the user accidently disconnects from the system (SIGHUP).

If an auto-save file exists and you is desire to recover data from the auto-save file, use the function recover_file. Whenever jed finds a file, it checks to see if an auto-save file exists as well as the file’s date. If the dates are such that the auto-save file is more recent jed will display a message in the mini-buffer alerting the user of this fact and that the function recover_file should be considered.

8 Status line and Windows

jed supports multiple windows. Each window may contain the same buffer or different buffers. A status line is displayed immediately below each window. The status line contains information such as the jed version number, the buffer name, “mode”, etc. Please beware of the following indicators:

**

buffer has been modified since last save.

%%

buffer is read only.

m

Mark set indicator. This means a region is being defined.

d

File changed on disk indicator. This indicates that the file associated with the buffer is newer than the buffer itself.

s

spot pushed indicator.

+

Undo is enabled for the buffer.

[Macro]

A macro is being defined.

[Narrow] Buffer is narrowed to a region of LINES.

9 Mini-Buffer

The Mini-Buffer consists of a single line located at the bottom of the screen. Much of the dialog between the user and jed takes place in this buffer. For example, when you search for a string, jed will prompt you for the string in the Mini-Buffer.

The Mini-Buffer also provides a direct link to the S-Lang interpreter. To access the interpreter, press Ctrl-X Esc and the S-Lang> prompt will appear in the Mini-Buffer. Enter any valid S-Lang expression for evaluation by the interpreter.

It is possible to recall data previously entered into the Mini-Buffer by using the up and down arrow keys. This makes it possible to use and edit previous expressions in a convenient and efficient manner.

9.1 Command Line Completion

The jed editor has several hundred built–in functions as well as many more written in the S-Lang extension language. Many of these functions are bound to keys and many are not. It is simply unreasonable to require the user to remember if a function is bound to a key or not and, if it is, to remember the key to which it is bound. This is especially true of those functions that are bound but rarely used. More often than not, one simply forgets the exact name or spelling of a function and requires a little help. For this reason, jed supports command line completion in the mini-buffer. This function, called emacs_escape_x, is bound to the key Esc X. This is one binding that must be remembered!

As an example, suppose that you are editing several buffers and you wish to insert the contents of one buffer into the current buffer. The function that does this is called insert_buffer and has no default key-binding. Pressing Esc X produces the prompt M-x. This prompt, borrowed from the Emacs editor, simply means that Esc X was pressed. Now type in and hit the space bar or the Tab key. In this context (completion context) the space bar and the Tab will expand the string in the Mini-Buffer up until it is no longer unique. In this case, insert_file and insert_buffer are only the two functions that start with in. Hence, in will expand to insert_at which point it becomes necessary to enter more information to uniquely specify the desired function. However, in a completion context, the space bar also has a special property that enables the user to cycle among the possible completions. For this example, hitting the space bar twice consecutively will produce the string insert_file and hitting it again produces the desired string insert_buffer.

The role of the space bar in completion is a point where Emacs and jed differ. Emacs will pop up a buffer of possible completions but jed expects the user to press the space bar to cycle among them. Both have there pros and cons. Frequently, one sees messages on the Usenet newsgroup gnu.emacs.help from Emacs users asking for the kind of completion jed employs.

9.2 File Names

jed takes every file name and “expands it” according to a set of rules which vary according to the Operating System. For concreteness, consider jed running under MS-DOS. Suppose the user reads a new file into the editor via the find_file command which emacs binds to Ctrl-X Ctrl-F. Then the following might be displayed in the mini-buffer:

      Find File: C:\JED\SLANG\

Here jed is prompting for a file name in the directory \JED\SLANG on disk C:. However, suppose the user wants to get the file C:\JED\SRC\VIDEO.C. Then the following responses produce equivalent filenames when jed expands them internally:

      Find File: C:\JED\src\video.c
      Find File: C:\JED\SLANG\..\src\video.c
      Find File: C:\JED\SLANG\../src/video.c

Note that the on MS-DOS systems, jed replaces the / with a \ and that case is not important. Now suppose you wish to get the file VIDEO.C from disk A:. The following are also valid:

      Find File: A:\video.c
      Find File: A:video.c
      Find File: C:\JED\SLANG\a:\video.c

In the last case, jed is smart enough to figure out what is really meant. Although the above examples are for MS-DOS systems, the rules also apply to Unix and VMS systems as well. The only change is the file name syntax. For example, on VMS

      sys$manager:[misc]dev$user:[davis.jed]vms.c
      dev$user:[davis.jed]vms.c

become equivalent filenames upon expansion. For unix, the following are equivalent:

      /user1/users/davis/jed/unix.c
      /usr/local/src//user1/users/davis/jed/unix.c
      /usr/local/src/~/jed/unix.c

Note the last example: the tilde character ~ always expands into the users HOME directory, in this case to

/user1/users/davis.

When jed writes a buffer out to a file, it usually prompts for a file name in the minibuffer displaying the directory associated with the current buffer. At this point a name can be appended to the directory string to form a valid file name or the user may simply hit the RET key. If the latter alternative is chosen, jed simply writes the buffer to the file already associated with the buffer. Once the buffer is written to a file, the buffer becomes attached to that file.

9.3 Buffer Name and File Name Completion

When jed prompts for a file name or a buffer name, the space bar and the Tab keys are special. Hitting the Tab key will complete the name that is currently in the minibuffer up until it is no longer unique. At that point, you can either enter more characters to complete the name or hit the space bar to cycle among the possible completions. The spacebar must be pressed at least twice to cycle among the completions.

On MSDOS and VMS, it is possible to use wildcard characters in the file name for completion purposes. For example, entering *.c and hitting the space bar will cycle among file names matching *.c. Unfortunately, this feature is not available on unix systems.

10 Basic Editing

Editing with jed is pretty easy— most keys simply insert themselves. Movement around the buffer is usually done using the arrow keys or page up and page down keys. If edt.sl is loaded, the keypads on VTxxx terminals function as well. Here, only the highlights are touched upon (cut/paste operations are not considered “highlights”). In the following, any character prefixed by the ^character denotes a Control character. On keyboards without an explicit Escape key, Ctrl-[ will most likely generate and Escape character.

A “prefix argument” to a command may be generated by first hitting the Esc key, then entering the number followed by pressing the desired key. Normally, the prefix argument is used simply for repetition. For example, to move to the right 40 characters, one would press Esc 4 0 followed immediately by the right arrow key. This illustrates the use of the repeat argument for repetition. However, the prefix argument may be used in other ways as well. For example, to begin defining a region, one would press the Ctrl-@ key. This sets the mark and begins highlighting. Pressing the Ctrl-@ key with a prefix argument will abort the act of defining the region and to pop the mark.

The following list of useful keybindings assumes that emacs.sl has been loaded.

Ctrl-L Redraw screen.

Ctrl-Undo (Control-underscore, also Ctrl-X u).

Esc q Reformat paragraph (wrap mode). Used with a prefix argument. will justify the paragraph as well.

Esc n narrow paragraph (wrap mode). Used with a prefix argument will justify the paragraph as well.

Esc ; Make Language comment (Fortran and C)

Esc \ Trim whitespace around point

Esc ! Execute shell command

Esc $ Ispell word (unix)

Ctrl-X ? Show line/column information.

‘ quoted_insert — insert next char as is (backquote key)

Esc s Center line.

Esc u Upcase word.

Esc d Downcase word.

Esc c Capitalize word.

Esc x Get M-x minibuffer prompt with command completion

Ctrl-X Ctrl-B pop up a list of buffers

Ctrl-X Ctrl-C exit jed

Ctrl-X 0 Delete Current Window

Ctrl-X 1 One Window.

Ctrl-X 2 Split Window.

Ctrl-X o Other window.

Ctrl-X b switch to buffer

Ctrl-X k kill buffer

Ctrl-X s save some buffers

Ctrl-X Esc Get S-Lang> prompt for interface to the S-Lang interpreter.

Esc . Find tag (unix ctags compatible)

Ctrl-@ Set Mark (Begin defining a region). Used with a prefix argument aborts the act of defining the region and pops the Mark.

10.1 Undo

One of jed’s nicest features is the ability to undo nearly any change that occurs within a buffer at the touch of a key. If you delete a word, you can undo it. If you delete 10 words in the middle of the buffer, move to the top of the buffer and randomly make changes, you can undo all of that too.

By default, the undo function is bound to the key Ctrl- (Ascii 31). Since some terminals are not capable of generating this character, it is also bound to the key sequence Ctrl-X u.

Due to the lack of virtual memory support on IBMPC systems, the undo function is not enabled on every buffer. In particular, it is not enabled for the *scratch* buffer. However, it is enabled for any buffer which is associated with a file. A “plus” character on the left hand side of the status line indicates that undo is enabled for the buffer. It is possible to enable undo for any buffer by using the toggle_undo function.

10.2 Marking Text (Point and Mark)

Many commands work on certain regions of text. A region is defined by the Point and the Mark The Point is the location of the current editing point or cursor position. The Mark is the location of a mark. The mark is set using the set_mark_cmd which is bound to Ctrl-@ (Control-2 or Control-Space on some keyboards). When the mark is set, the m mark indicator will appear on the status line. This indicates that a region is being defined. Moving the cursor (Point) defines the other end of a region. If the variable HIGHLIGHT is non-zero, jed will highlight the region as it is defined.

Even without highlighting, it is easy to see where the location of the mark is by using the exchange command which is bound to Ctrl-X Ctrl-X. This simply exchanges the Point and the Mark. The region is still intact since it is defined only by the Point and Mark. Pressing Ctrl-X Ctrl-X again restores the mark and Point back to their original locations. Try it.

10.3 Tab Issues.

Strictly speaking, jed uses only fixed column tabs whose size is determined by the value of the TAB variable. Setting the TAB variable to 0 causes jed to not use tabs as whitespace and to display tabs as Ctrl-I. Please note that changing the tab settings on the terminal will have no effect as far as jed is concerned. The TAB variable is local to each buffer allowing every buffer to have its own tab setting. The variable TAB_DEFAULT is the tab setting that is given to all newly created buffers. The default value for this variable is 8 which corresponds to eight column tabs.

jed is also able to “simulate” arbitrary tabs as well through the use of user defined tab stops. Calling the function edit_tab_stops allows the user to interactively set the tab stops. That is, one simply presses Esc X to get the M-x prompt and enters edit_tab_stops. A window will pop open displaying the current tab settings. To add a tab stop, simply place a T in the appropriate column. Use the space bar to remove a tab stop.

Here an argument is presented in favor of simulated tabs over real tab stops. First, consider what a “tab” really is. A “tab” in a file is nothing more than a character whose ASCII value is 9. For this reason, one also denotes a tab as ^I (Ctrl-I). Unlike most other ASCII characters, the effect of the tab character is device dependent and is controlled through the device tab settings. Hence, a file which displays one way on one device may look totally different on another device if the tab settings do not correspond. For this reason, many people avoid tabs altogether and others the adopt “standard” of eight column tabs. Even though people always argue about what the correct tab settings should be, it must be kept in mind that this is primarily a human issue and not a machine issue.

On a device employing tab stops, a tab will cause the cursor to jump to the position of the next tab stop. Now consider the effect of changing the tab settings. Assume that in one part of a document, text was entered using the first setting and in another part, the second setting was used. When moving from the part of the document where the current tab setting is appropriate to the part where the other tab setting was used will cause the document to look unformatted unless the appropriate tab settings are restored. Wordprocessors store the tab settings in the file with the text so that the tabs may be dynamically changed to eliminate such unwanted behavior. However, text editors such as jed, vi, Emacs, EDT, EVE (TPU), etc, do not store this information in the file. jed avoids this problem by using simulated tabs. When using simulated tabs, tabs are not really used at all. Rather jed inserts the appropriate number of spaces to achieve the desired effect. This also has the advantage of one being able to cut and paste from the part of a document using one tab setting to another part with a different tab setting. This simple operation may lead to unwanted results on some wordprocessors as well as those text editors using real tab stops.

10.4 Searching

jed currently has two kinds of searches: ordinary searches and incremental searches. Both types of searches have forward and backward versions. The actual functions for binding purposes are:

      search_forward
      search_backward
      isearch_forward
      isearch_backward

There is also the occur function which finds all occurrences of a single word (string). This function has no backwards version. By default it is not bound to any keys, so to use it, occur must be entered at the M-x prompt (Esc X) or one is always free to bind it to a key.

In the following only the incremental search is discussed.

The default type of search in Emacs in the incremental search. However, since this type of search is confusing to the uninitiated, the ordinary type of search has been made the default in jed’s Emacs emulation. For the traditional emacs keybinding, it is up to the user to provide the keybinding.

As the name suggests, an incremental search performs a search incrementally. That is, as you enter the search string, the editor begins searching right away. For example, suppose you wish to search for the string apple. As soon as the letter a is entered into the incremental search prompt, jed will search for the first occurrence of a. Then as soon as the p is entered, jed will search from the current point for the string ap, etc. This way, one is able to quickly locate the desired string with only a minimal amount of information.

Unlike the “ordinary” search, the incremental search is not terminated with the Enter key. Hitting the Enter key causes jed to search for the next occurrence of the string based on the data currently entered at the prompt. The search is terminated with the Esc key.

Finally, the DEL key (Ctrl-?) is used to erase the last character entered at the search prompt. In addition to erasing the last character of the search string, jed will return back to the location of the previous match. Erasing all characters will cause the editor to return to the place where the search began. Like many things, this is one of those that is easier to do than explain. Feel free to play around with it.

10.5 Rectangles

jed has built-in support for the editing of rectangular regions of text. One corner of rectangle is defined by setting the mark somewhere in the text. The Point (cursor location) defines the opposite corner of the rectangle.

Once a rectangle is defined, one may use the following functions:

kill_rect

Delete text inside the rectangle saving the rectangle in the internal rectangle buffer.

n_rect

Push all text in the rectangle to the right outside the rectangle.

copy_rect

Copy text inside the rectangle to the internal rectangle buffer.

blank_rect

Replace all text inside the rectangle by spaces.

The function insert_rect inserts a previously killed or copied rectangle into the text at the Point.

These functions have no default binding and must be entered into the MiniBuffer by pressing Esc X to produce the M-x prompt.

10.6 Sorting

jed is capable of sorting a region of lines using the heapsort algorithm. The region is sorted alphabetically based upon the ASCII values of the characters located within a user defined rectangle in the region. That is, the rectangle simply defines the characters upon what the sort is based. Simply move to the top line of the region and set the mark on the top left corner of the rectangle. Move to the bottom line and place the point at the position which defines the lower right corner of the rectangle. Press Esc X to get the M-x prompt and enter sort As as example, consider the following data:

Fruit:             Quantity:
lemons                3
pears                37
peaches             175
apples              200
oranges              56

To sort the data based upon the name, move the Point to the top left corner of the sorting rectangle. In this case, the Point should be moved to the l in the word lemons. Set the mark. Now move to the lower right corner of the rectangle which is immediately after the s in oranges. Pressing Esc X and entering sort yields:

Fruit:             Quantity:
apples              200
lemons                3
oranges              56
peaches             175
pears                37

Suppose that it is desired to sort by quantity instead. Looking at the original (unsorted) data, move the Point to two spaces before the 3 on the line containing lemons. The cursor should be right under the u in Quantity. Set the mark. Now move the Point to immediately after 56 on the oranges line and again press Esc X and enter sort. This yields the desired sort:

Fruit:             Quantity:
lemons                3
pears                37
oranges              56
peaches             175
apples              200

11 Modes

jed supports two internal modes as well as user defined modes. The two internal modes consist of a “C” mode for C Language programming and a “Wrap” mode for ordinary text editing. Examples of user defined modes are Fortran mode and DCL mode.

Online documentation is provided for nearly every mode jed defines. For help on the current mode, press Esc X and enter describe_mode. A window will appear with a short description of the special features of the mode as well as a description of the variables affecting the mode.

11.1 Wrap Mode

In this mode, text is wrapped at the column given by the WRAP variable. The default is 78. The text does not wrap until the cursor goes beyond the wrap column and a space is inserted.

11.1.1 Formatting paragraphs

Paragraph delimiters are: blank lines, lines that begin with either a percent character, %, or a backslash character \. This definition is ideally suited for editing LaTEX documents. However, it is possible for the user to change this definition. See the discussion of the hook, is_paragraph_separator, in the section on hooks for explicit details on how to do this.

The paragraph is formatted according to the indentation of the current line. If the current line is indented, the paragraph will be given the same indentation. The default binding for this function is Esc q.

In addition, a paragraph may be “narrowed” by the narrow_paragraph function which is bound to Esc N by default. This differs from the ordinary format_paragraph function described above in that the right margin is reduced by an amount equal to the indentation of the current line. For example:

      This paragraph is the result of using the
      function ``narrow_paragraph''.  Note how the
      right margin is less here than in the above
      paragraph.

Finally, if either of these functions is called from the keyboard with a prefix argument, the paragraph will be justified as well. For example, pressing Esc 1 Esc N on the previous paragraph yields:

      This paragraph  is  the  result   of   using  the
      function   ``narrow_paragraph''.    Note  how the
      right margin   is less here  than  in   the above
      paragraph.

See the discussion of format_paragraph_hook in the section on hooks for details on how this is implemented.

11.2 Smart Quotes

You have probably noticed that many key words in this document are quoted in double quotes like “this is double quoted” and ‘this is single quoted’. By default, the double quote key (”) and single quote key (’) are bound to the function text_smart_quote. With this binding and in wrap mode, the single quote key inserts a single quote with the “proper” orientation and the double quote key inserts two single quotes of the “proper” direction. To turn this off, rebind the keys to self_insert_cmd. Some modes already do this (e.g., EDT).

This brings up the question: if the double quote key is bound to text_smart_quote then how does one insert the character (”)? The most common way is to use the quoted_insert function which, by default, is bound to the single backquote (‘) key. This is the same mechanism that is used to insert control characters. The other method is to use the fact that if the preceding character is a backslash, \, the character simply self inserts. Again, this is ideal for writing TEX documents.

11.3 C Mode

C Mode facilitates the editing of C files. Much of the latter part of the development of the jed editor was done using this mode. This mode may be customized by a judicious choice of the variables C_INDENT and C_BRACE as well as the bindings of the curly brace keys { and }. Experiment to find what you like or write your own using the S-Lang interface.

By default, the Enter key is bound to the function newline_and_indent. This does what its name suggests: inserts a newline and indents. Again, some modes may rebind this key. In addition, the keys {, }, and Tab are also special in this mode. The Tab key indents the current line and the { and } keys insert themselves and reindent. If you do not like any of these bindings, simply rebind the offending one

to self_insert_cmd.

Finally, the key sequence Esc ; is bound to a function called c_make_comment. This function makes and indents a C comment to the column specified by the value of the variable C_Comment_Column. If a comment is already present on the line, it is indented.

11.4 Fortran Mode

Fortran Mode is written entirely in S-Lang and is designed to facilitate the writing of Fortran programs. It features automatic indentation of Fortran code as well as automatic placement of Fortran statement Labels.

In this mode, the keys 0-9 are bound to a function for_elebel which does the following:

1. Inserts the calling character (0-9) into the buffer.

2. If the character is preceded by only other digit characters, it assumes the character is for a label and moves it to the appropriate position.

3. Reindents the line.

This function is very similar to the one Emacs uses for labels.

12 Keyboard Macros

jed is able to record a series of keystrokes from the terminal and replay them. The saved series of keystrokes is known as a keyboard macro. To begin a keyboard macro, simply enter the begin keyboard macro key sequence which is bound to Ctrl-X ( if emacs.sl is loaded. To stop recording the keystrokes, enter Ctrl-X ). Then to “execute” the macro, press Ctrl-X e. Please note that it is illegal to execute a macro while defining one and doing so generates an error. A macro can be aborted at anytime by pressing the Ctrl-G key.

One nice feature jed includes is the macro_query function. That is, while defining a macro, the key sequence Ctrl-X q will cause jed to issue the prompt Enter String: in the minibuffer. Any string that is entered will be inserted into the buffer and the process of defining the macro continues. Every time the macro is executed, jed will prompt for a NEW string to be inserted.

Any time an error is generated, the process of defining the macro is aborted as well as execution of the macro. This is very useful and may be exploited often. For example, suppose you want to trim excess whitespace from the end of ALL lines in a buffer. Let us also suppose that the number of lines in the buffer is less than 32000. Then consider the following keystrokes:

      Ctrl-X (                      (begin macro)
      Ctrl-E                        (goto end of line)
      ESC                           (trim whitespace)
      Down Arrow                    (go down one line)
      Ctrl-X )                      (end macro)

Now the macro has been defined. So move to the top of the buffer and execute it 32000 times:

      ESC <                         (top of buffer)
      ESC 3 2 0 0 0                 (repeat next command 32000 times
      Ctrl-X e                      (execute macro)

If the buffer has less than 32000 lines, the end of the buffer will be reached and an error will be generated aborting the execution of the macro.

13 Shells and Shell Commands

The default binding to execute a shell command and pump the output to a buffer is Esc !. jed will prompt for a command line and spawn a subprocess for its execution.

Strictly speaking, jed does not support interactive subprocesses. However, jed includes S-Lang code that “emulates” such a subprocess. It may invoked by typing shell at the M-x minibuffer prompt. A window will be created with a buffer named *shell* attached to it. Any text entered at the system dependent shell prompt will be executed in a subprocess and the result stuffed back in the shell buffer. Don’t try to execute any commands which try to take over the keyboard or the screen or something undesirable may happen. Examples of types of stupid commands are spawning other editors, logging in to remote systems, et cetera. Even chdir is stupid since its effect is not permanent. That is,

      > cd ..
      > dir

will not do what might naively be expected. That is, the two commands above are not equivalent to the single command dir ...

14 Getting Help

jed’s help functions are bound to Ctrl-H by default. For example, Ctrl-H C will show what function a key carries out, Ctrl-H i will run jed’s info reader, Ctrl-H f will give help on a particular S-Lang function, etc. However, some modes may use the Ctrl-H key for something else. For example, if EDT mode is in effect, then Ctrl-H may be bound to bol which causes the cursor to move to the beginning of the line. See the section on EDT for more information.

If jed is properly installed, this entire document is accessable from within the editor using jed’s info reader. Ctrl-H i will load info_mode allowing the user to browse the document as well as other “info” documents.

15 Editing Binary Files

jed may edit binary files as long as the proper precautions are taken. On IBMPC systems, this involves calling the S-Lang function set_file_translation with an integer argument. If the argument is 0, files are opened as text files; otherwise, they are opened in binary mode. There is no need to call this function for other systems. However, beware of the user variable ADD_NEWLINE which if non zero, a newline character will be appended to the file if the last character is not a newline character. If you are going to edit binary files, it is probably a good idea to set this variable to zero.

16 Dired— the Directory editor

In addition to editing files, jed is also able to rename and delete them as well. jed’s Dired mode allows one to do just this is a simple and safe manner.

To run dired, simply press Esc X and enter dired at the prompt. jed will load dired.sl and prompt for a directory name. Once the directory is given, jed will display a list files in the directory in a buffer named *dired*. One may use normal buffer movement keys to move around this buffer. To delete one or more files, use the d key to “tag” the files. This in itself does not delete them; rather, it simply marks them for deleting. A capital ‘D’ will appear in the left margin to indicate that a file has been tagged. Simply hit the u key to untag a file. The delete key will also untag the previously tagged file.

To actually delete the tagged files, press the ‘x’ key. This action causes jed to display a list of the tagged files in a separate window and prompt the user for confirmation. Only when the proper confirmation is given, will the file be deleted.

Renaming a file is just as simple. Simply move to the line containg the name of the file that you wish to rename and hit the ‘r’ key. jed will prompt for a filename or a directory name. If a directory is given, the file will be moved to the new directory but will keep the name. However, for the operation to succeed, the file must be one the same file system. To rename tagged files to a different directory residing on the same file system, use the m key. This has the effect of moving the tagged file out of the current directory to the new one.

One may also use the f key to read the file indicated by the cursor position into a buffer for editing. If the file is a directory, the directory will be used for dired operations. In addition, one may also use the v to simply “view” a file.

Finally, the g key will re-read the current directory and the h and ? keys provide some help.

17 Mail

This section applies to Unix and VMS systems only. On these systems, it is possible to compose and send mail directly using jed. This assumes that the Unix system has /usr/ucb/mail. It is trivial to modify mail.sl to support another Unix mailer. For VMS, jed uses the callable mail interface present on VMS versions 5.0 and later.

The default binding for the mail is Ctrl-X m. Alternatively, one may press Esc X and enter mail at the M-x prompt. The mail function will cause a window to open with a buffer called *mail* which contains the three lines:

      To:
      Subject:
      ---text follows this line---

Simply enter the email address of the person that you want to send the mail to on the line containing To: and put the subject of the message on the next line labeled Subject:. The text that you wish to mail follows the line labeled ---text follows this line--- which is used by jed as a marker. After you have composed the mail message, press Esc X and enter send at the M-x prompt. For example, the following is an email requesting to be put on the jed mailing list:

      To: davis@space.mit.edu
      Subject: jed mailing list
      ---text follows this line---
      Hi,
         Please add me to the JED mailing list so that I may be notified
      of upcoming releases of JED.
      --Maria

For VMS systems, the above example will probably fail because an internet address has been used for the example. For systems using a TCP/IP package, it may be necessary to change davis@space.mit.edu to

something like smtp%"davis@space.mit.edu".

The mail function looks for a user defined hook called mail_hook and execute it if it exists. This hook may be used to bind certain keys in the keymap associated with the *mail* buffer. For example,

      define mail_hook ()
      {
        local_unsetkey ("^C");
        local_setkey ("send", "^C^C");
      }

defines the key Ctrl-C Ctrl-C in the mail keymap to perform the send function. Other possibilities include binding a key sequence, say Ctrl-C Ctrl-W, to a function that inserts the contents of a signature file.

18 Customization

To extend jed, it is necessary to become familiar with the S-Lang programming language. S-Lang not a standalone programming language like C, Pascal, etc. Rather it is meant to be embedded into a C program. The S-Lang programming language itself provides only arithmetic, looping, and branching constructs. In addition, it defines a few other primitive operations on its data structures. It is up to the application to define other built-in operations tailored to the application. That is what has been done for the jed editor. See the document slang.txt for S-Lang basics as well as the jed Programmer’s Manual for functions jed has added to the language. In any case, look at the *.sl files for explicit examples.

For the most part, the average user will simply want to rebind some keys and change some variables (e.g., tab width). Here I discuss setting keys and the predefined global variables.

18.1 Setting Keys

Defining a key to invoke a certain function is accomplished using the setkey function. This function takes two arguments: the function to be executed and the key binding. For example, suppose that you want to bind the key Ctrl-A to cause the cursor to go to the beginning of the current line. The jed function that causes this is bol (See the jed Programmer’s Manual for a complete list of functions). Putting the line:

      setkey ("bol", "^A");

in the startup file jed.rc (.jedrc) file will perform the binding. Here ^A consists of the two characters ^and A which jed will interpret as the single character Ctrl-A. For more examples, see either of the

S-Lang files emacs.sl or edt.sl.

The first argument to the setkey function may be any S-Langexpression. Well, almost any. The only restriction is that the newline character cannot appear in the expression. For example, the line

      setkey ("bol();skip_white ();",  "^A");

defines the Ctrl-A key such that when it is pressed, the editing point will move the beginning of the line and then skip whitespace up to the first non-whitespace character on the line.

In addition to being able to define keys to execute functions, it is also possible to define a key to directly insert a string of characters. For example, suppose that you want to define a key to insert the string int main(int argc, char **argv) whenever you press the key Esc m. This may be accomplished as follows:

      setkey (" int main(int argc, char **argv)", "\em");

Notice two things. First of all, the key sequence Esc m has been written as "\em" where \e will be interpreted by jed as Esc. The other salient feature is that the first argument to setkey, the “function” argument, begins with a space. This tells jed that it is not be interpreted as the name of a function; rather, the characters following the space are to be inserted into the buffer. Omitting the space character would cause jed to execute a function called int main(int argc, char **argv) which would fail and generate an error.

Finally, it is possible to define a key to execute a series of keystrokes similar to a keyboard macro. This is done by prefixing the “function” name with the @ character. This instructs jed to interpret the characters following the @ character as characters entered from the keyboard and execute any function that they are bound to. For example, consider the following key definition which will generate a C language comment to comment out the current line of text. In C, this may be achieved by inserting symbol "/*" at the beginning of the line and inserting "*/" at the end of the line. Hence, the sequence is clear (Emacs keybindings):

1. Goto the beginning of the line: Ctrl-A or decimal "\001".

2. Insert /*.

3. Goto end of the line: Ctrl-E or decimal \005.

4. Insert */

To bind this sequence of steps to the key sequence Esc ;, simply use

     setkey("@\001/*\005*/", "\e;");

Again, the prefix @ lets jed know that the remaining characters will carry out the functions they are currently bound to. Also pay particular attention to the way Ctrl-A and Ctrl-E have been written. Do not attempt to use the ^to represent “Ctrl”. It does not have the same meaning in the first argument to the setkey function as it does in the second argument. To have control characters in the first argument, you must enter them as \xyz where xyz is a three digit decimal number coinciding with the ASCII value of the character. In this notation, the Esc character could have been written as \027. See the S-Lang Programmer’s Reference Manual for further discussion of this notation.

The setkey function sets a key in the global keymap from which all others are derived. It is also possible to use the function local_setkey which operates only upon the current keymap which may or may not be the global map.

18.2 Predefined Variables

jed includes some predefined variables which the user may change. By convention, predefined variables are in uppercase. The variables which effect all modes include:

BLINK

(1) if non-zero, blink matching parenthesis.

TAB_DEFAULT

(8) sets default tab setting for newly created buffers to specified number of columns.

TAB

Value of tab setting for current buffer.

ADD_NEWLINE

(1) adds newline to end of file if needed when writing it out to the disk.

META_CHAR

(-1) prefix for chars with high bit set (see section on eight bit clean issues for details)

DISPLAY_EIGHT_BIT

see section on eight bit clean issues.

COLOR

(23) IBMPC background color (see jed.rc for meaning)

LINENUMBERS

(0) if 1, show current line number on status line

WANT_EOB

(0) if 1, [EOB] denotes end of buffer.

TERM_CANNOT_INSERT

(0) if 1, do not put the terminal in insert mode when writing to the screen.

IGNORE_BEEP

(0) do not beep the terminal when signalling errors

In addition to the above, there are variables which affect only certain modes. See the section on modes for details.

18.3 Hooks

A hook is a user defined function that jed calls under certain conditions which allow the user to modify default actions. For example, when jed starts up it looks for the existence of a user defined function command_line_hook. If this function exists, jed calls the function. What the function does is completely arbitrary and is left to the discretion of the user. The startup file, site.sl, defines such a function which reads in the files listed on the command line. It is also this function which loads the jed.rc startup file. Unlike the other hooks, this one must be present in the file site.sl since it is the only file loaded before calling the hook.

After the startup files are loaded, jed calls the hook jed_startup_hook immediately before entering the main editor loop. This hook is useful to modify certain data structures which may not have existed when the startup files were loaded.

In addition to the above hooks, jed currently also looks for:

suspend_hook

function to be executed before suspending

resume_hook

function that gets carried out after suspension

exit_hook

gets executed before exiting jed

mode_hook

sets buffer mode based on filename extension

find_file_hook

called before file is read into a buffer. It currently checks for presence of autosave file and warns user if it is more recent than file.

See site.sl for explicit examples of the above hooks.

Another useful hook is is_paragraph_separator. This hook is called when jed searches for the beginning or end of a paragraph. This search is performed by all paragraph formatting functions as well as the forward and backward paragraph movement commands. As jed performs the search, it moves from one line to another testing the line to see if it separates a paragraph. The function of the hook is to make this decision and return zero if the line does not separate paragraphs or return one if it does. The default value of this hook may be written in S-Lang as

     define is_paragraph_separator ()
     {
       bol ();
       if (looking_at ("\\")) return 1;
       if (looking_at ("%")) return 1;
       skip_white(); eolp ();
     }

A related hook called after a paragraph is formatted is format_paragraph_hook. This hook is only called if either format_paragraph or narrow_paragraph is called with a prefix digit argument. For example, format_paragraph is bound to Esc q. Simply pressing this key sequence will call format_paragraph but format_paragraph_hook will not be called. However, pressing Esc 1 followed by Esc q will result in a call to format_paragraph_hook. Currently, this hook simply justifies the paragraph. That is, it fills each line in the paragraph such that the line ends at the right margin, which is defined by the WRAP variable.

18.4 S-Lang Programming Hints (Debugging)

This section assumes some knowledge about S-Lang and is designed to explain how to debug S-Lang routines quickly. For information about S-Lang, read slang.txt.

There are two ways of loading a file of S-Lang code into jed. The most common way is through the function evalfile. If an error occurs while loading a file, jed will give some indication of where the problem lies by displaying the line number and the offending bit of S-Lang code in the minibuffer. In practice though, this can be quite inefficient. The evalfile function is primarily designed to load debugged and tested S-Lang code.

The best way to develop and test S-Lang code with jed is to use the function evalbuffer. Simply load the piece of code into jed as an ordinary file, press Esc X and enter the function evalbuffer If the piece of code in the buffer has any syntax errors, jed will put the cursor on the error. This is the best way to spot compile time errors such as syntax errors. However, this will not catch runtime errors.

When a runtime error occurs, jed will put the cursor on the top level function where the original call was made and NOT the actual location of the function. To aid in determining where an error occurs, jed can be made to give a symbolic traceback. As the S-Lang runtime stack unwinds, S-Lang will simply print the name of function at that particular level. If the function includes local variables, their values will be dumped as well. Hence, it is easy to quickly narrow the location of an error down to function where the error occurs. By default, the traceback is disabled. The traceback is enabled by setting the S-Lang variable _traceback to a non-zero value. It is simpliest to just press Ctrl-X Esc and enter _traceback = 1 at the S-Lang prompt. This is one of those times where one needs access to the S-Lang> prompt and not the M-x prompt. For example, consider the following piece of code:

      define fun_two () {forever {}}   % loops forever
      define fun_one () {fun_two ()}   % calls fun_two-- never returns

Simply enter the above into an empty jed *scratch* buffer, then press Ctrl-X Esc and enter:

      _traceback = 1; () = evalbuffer (); fun_one ();

This will turn on tracebacks, evaluate the buffer and call the function fun_one. jed will then be put into an infinite loop which can only be stopped by pressing the abort character which by default is Ctrl-G. Doing so, will produce the traceback messages

       S-Lang Traceback: fun_two
       S-Lang Traceback: fun_one

in addition to the error message User Break!. Of course, this technique only narrows down the source of an error to a particular function. To proceed further, it may necessary to put “print” statements at suitable places in the function. There are several ways to do this:

1. Use the insert function to insert the contents of a variable into the current buffer.

2. Use the error function to abort the function and display the value of a variable in the minibuffer.

3. Use the message function to display the value of a variable in the minibuffer. Unlike error, the message function does not abort the execution of the function.

Since each of these functions require a string argument, it is usually best to call the string function first for the conversion followed by the output function. This has to be done anyway if it is desired to get the contents of an integer variable. Although the second approach is prehaps the most useful in practice, it is somtimes appropriate to use a combination of these techniques.

Finally, to print the entire stack, one can use the print_stack function. This function dumps the S-Lang runtime stack into the *traceback* buffer.

Since S-Lang is an interpreted language, judicious application of the above techniques should lead very quickly to the source of any errors.

19 Eight Bit Clean Issues

19.1 Displaying Characters with the High Bit Set

There are several issues to consider here. The most important issue is how to get jed to display 8 bit characters in a “clean” way. By “clean” I mean any character with the high bit set is sent to the display device as is. This is achieved by putting the line:

      DISPLAY_EIGHT_BIT = 1;

in the jed.rc (.jedrc) startup file. European systems might want to put this in the file site.sl for all users. The default is 1 so unless its value has been changed, this step may not be necessary.

There is another issue. Suppose you want to display 8 bit characters with extended Ascii codes greater than or equal to some value, say 160. This is done by putting DISPLAY_EIGHT_BIT = 160;. I believe that ISO Latin character sets assume this. This is the default value for Unix and VMS systems.

19.2 Inputting Characters with the hight bit Set

Inputting characters with the high bit set into jed is another issue. How jed interprets this bit is controlled by the variable META_CHAR. What happens is this: When jed reads a character from the input device with the high bit set, it:

1. Checks the value of META_CHAR. If this value is -1, jed simply inserts the character into the buffer.

2. For any other value of META_CHAR in the range 0 to 255, jed returns two 7-bit characters. The first character returned is META_CHAR itself. The next character returned is the original character but with the high bit stripped.

The default value of META_CHAR is -1 which means that when jed sees a character with the high bit set, jed leaves it as is. Please note that a character with the high bit set it cannot be the prefix character of a keymap. It can be a part of the keymap but not the prefix.

Some systems only handle 7-bit character sequences and as a result, jed will only see 7-bit characters. jed is still able to insert any character in the range 0-255 on a 7-bit system. This is done through the use of the quoted_insert function which, by default, is bound to the backquote key `. If the quoted_insert function is called with a digit argument (repeat argument), the character with the value of the argument is inserted into the buffer. Operationally, one hits Esc, enters the extended Ascii code and hits the backquote key. For example, to insert character 255 into the buffer, simply press the following five keys: Esc 2 5 5 ‘.

19.3 Upper Case - Lower Case Conversions

The above discussion centers around input and output of characters with the high bit set. How jed treats them internally is another issue and new questions arise. For example, what is the uppercase equivalent of a character with ASCII code 231? This may vary from language to language. Some languages even have characters whose uppercase equivalent correspond to multiple characters. For jed, the following assumptions have been made:

• Each character is only 8 bits.

• Each character has a unique uppercase equivalent.

• Each character has a unique lowercase equivalent.

It would be nice if a fourth assumption could be made:

• The value of the lowercase of a character is greater than or equal to its uppercase counterpart.

However, apparently this is not possible since most IBMPC character sets violate this assumption. Hence, jed does not assume it. Suppose X is the upper case value of some character and suppose Y is its lower case value. Then to make jed aware of this fact and use it case conversions, it may be necessary to put a statement of the form:

     define_case (X, Y);

in the startup file. For example, suppose 211 is the uppercase of 244. Then, the line

      define_case (211, 244);

will make jed use this fact in operations involving the case of a character.

This has already been done for the ISO Latin 1 character set. See the file iso-latin.sl for details. For MSDOS, this will not work. Instead use the files dos437.sl and dos850.sl. By default, jed’s internal lookup tables are initialized to the ISO Latin set for Unix and VMS systems and to the DOS 437 code page for the IBMPC. To change the defaults, it is only necessary to load the appropriate file. For example, to load dos850.sl definitions, put

      evalfile ("dos850"); pop ();

in the startup file (e.g., site.sl). In addition to uppercase/lowercase information, these files also contain word definitions, i.e., which characters constitute a “word”.

20 Miscellaneous

20.1 Abort Character

The abort character (Ctrl-G by default) is special and should not be rebound. On the IBMPC, the keyboard interrupt 0x09 is hooked and a quit condition is signaled when it is pressed. For this reason, it should not be used in any keybindings. A similar statement holds for the other systems.

This character may be changed using the function set_abort_char Using this function affects all keymaps. For example, putting the line

      set_abort_char (30);

in your jed.rc file will change the abort character from its current value to 30 which is Ctrl-^.

20.2 Input Translation

By using the function map_input the user is able to remap characters input from the terminal before jed’s keymap routines have a chance to act upon them. This is useful when it is difficult to get jed to see certain characters. For example, consider the Ctrl-S character. This character is especially notorious because many systems use it and Ctrl-Q for flow control. Nevertheless Emacs uses Ctrl-S for searching. Short of rebinding all keys which involve a Ctrl-S how does one work with functions that are bound to key sequences using Ctrl-S? This is where map_input comes into play. The map_input function requires two integer arguments which define how a given ascii character is to be mapped. Suppose that you wish to substitute Ctrl-\ for Ctrl-S everywhere. The line

      map_input (28, 19);

will do the trick. Here 28 is the ascii character of Ctrl-\ and 19 is the ascii character for the Ctrl-S.

As another example, consider the case where the backspace key sends out a Ctrl-H instead of the DEL character (Ctrl-?).

      map_input (8, 127);

will map the Ctrl-H (ascii 8) to the delete character (ascii 127).

20.3 Display Sizes

On VMS and unix systems, the screen size may be changed to either 80 or 132 columns by using the functions w80 and w132 respectively. Simply enter the appropriate function name at the M-x prompt in the minibuffer. The default binding for access to the minibuffer is Esc X. Most window systems, e.g., DECWindows, allow the window size to be changed. When this is done, jed should automatically adapt to the new size.

On the PC, at this time the screen size cannot be changed while jed is running. Instead it is necessary to exit jed first then set the display size and rerun jed.